# Planning In Natural Language Improves LLM Search For Code Generation

### v1

@@ Abstract
While scaling training compute has led to remarkable improvements in large language models (LLMs), scaling inference compute has not yet yielded analogous gains. We hypothesize that a core missing component is a lack of diverse LLM outputs, leading to inefficient search due to models repeatedly sampling highly similar, yet incorrect generations. We empirically demonstrate that this lack of diversity can be mitigated by searching over candidate plans for solving a problem in natural language. Based on this insight, we propose PlanSearch, a novel search algorithm which shows strong results across HumanEval+, MBPP+, and LiveCodeBench (a contamination-free benchmark for competitive coding). PlanSearch generates a diverse set of observations about the problem and then uses these observations to construct plans for solving the problem. By searching over plans in natural language rather than directly over code solutions, PlanSearch explores a significantly more diverse range of potential solutions compared to baseline search methods. Using PlanSearch on top of Claude 3.5 Sonnet achieves a state-of-the-art pass@200 of 77.0% on LiveCodeBench, outperforming both the best score achieved without search (pass@1 = 41.4%) and using standard repeated sampling (pass@200 = 60.6%). Finally, we show that, across all models, search algorithms, and benchmarks analyzed, we can accurately predict performance gains due to search as a direct function of the diversity over generated ideas.
@@ 1 Introduction
The bitter lesson [40] famously posits that two forms of scaling trump everything else: learning and search. While recent advances in large language models have removed all doubt on the effectiveness of learning, search has not yet proven its value for large language models, despite its success with classical machine learning techniques [10, 37, 38, 7, 8, 4, 17].
@@
Furthermore, the lack of diversity is particularly harmful for search algorithms. In the most egregious of cases with little to no diversity, such as greedy decoding, repeated sampling from the model returns highly similar programs, resulting in minimal gain from additional inference-time compute. This diversity problem is also not reflected in many public leaderboards (e.g. LMSYS Chatbot Arena [14], LiveCodeBench [22], OpenLLMLeaderboard [1]), which often report only the pass rate from a single sample of the model, ignoring an entire dimension along which to compare models. While the performance of one sample is the primary metric of relevance for applications such as chatbots, as users typically are sensitive to latency, this single scalar is insufficient to fully capture the quality of a model when it is allowed to use more inference time compute.
@@
In this paper, we explore several directions for improving the diversity of LLMs at inference time. We hypothesize that the right axis of diversity to search over is the natural language conceptual/idea space, and we validate our hypothesis across several experiments. First, we show that models can produce the correct final program when fed the correct solution sketches, where these sketches have been “backtranslated” from passing solution code into sketches in idea space (Section 3.2). Second, we show that when models are asked to generate their own ideas before implementing them on LiveCodeBench (IdeaSearch), their accuracy conditioned on a particular sketch trends towards either 0% or 100%, suggesting that most of the variance in passing a particular problem is captured by whether the sketch is correct rather than any other factor. These two experiments suggest a natural method to improving LLM search for code generation: by searching for the correct idea to implement.
Guided by this principle of maximizing exploration of ideas, we propose PlanSearch. In contrast to many existing search methods that search over individual tokens [44, 46], lines of code [24], or even entire programs [25], PlanSearch searches over possible plans for solving the problem at hand, where a plan is defined as a collection of high level observations and sketches helpful to solve a particular problem (Figure 2). To generate novel plans, PlanSearch generates a number of observations about the problem, before combining these observations into a candidate plan for solving the problem. This is done for every possible subset of the generated observations to maximally encourage exploration in idea space, before the codes are eventually all translated into a final code solution (Section 4.2). We find that searching over plans outperforms both standard repeated sampling and directly searching over ideas (IdeaSearch, introduced in Section 4.1.2) in terms of effectively using compute at inference time.
Applying PlanSearch on top of Claude 3.5 Sonnet achieves a state-of-the-art pass@200 of 77.0% on LiveCodeBench, outperforming both the best score achieved without search (pass@1 = 41.4%) and the standard best-of-n sampling score (pass@200 = 60.6%). Furthermore, consistent with recent findings on the effectiveness of search on top of small models [12, 6, 5, 42], allowing PlanSearch based on a small model (GPT-4o-mini) outperforms larger models not augmented with search after merely 4 attempts. Evaluations of PlanSearch across two other coding benchmarks, HumanEval+ and MBPP+ [26], suggest similar improvements.
@@ 2.1 Search in Classical AI
Classical search algorithms like breadth-first search, depth-first search, and A* search have been widely used for pathfinding, planning, and optimization [34]. More advanced search techniques like Monte Carlo tree search (MCTS) have achieved remarkable success in domains like game playing, enabling superhuman performance in Go [37, 38], poker [7, 8] and Diplomacy [17]. More recently, Jones [23] find scaling laws for the performance of AI systems in board games, where ELO improves logarithmically with the amount of compute spent at inference.
@@ 2.2 Search with Language Models
Regarding searching over plans in natural language, several approaches have also proposed generalizing chain-of-thought [41] reasoning into a search-like process, such as Tree of Thoughts [43] and Reasoning via Planning [20]. However, prior methods have largely demonstrated effectiveness on somewhat contrived problems designed to highlight the power of search, such as the game of 24, or classic planning benchmarks such as Blocksworld [27], where both benchmarks are easier to solve by explicitly considering many options, and where the ‘steps’ over which to search over are fairly obvious. By contrast, most real-world planning is used to assist in domains that are complex enough to benefit from, but not require, the additional exploration of plans. We demonstrate that PlanSearch, which plans in natural language, outperforms baseline search methods in one such domain: code generation. Moreover, our analysis reveals the underlying reason that such search is effective: it increases the diversity of the generated ideas, allowing more efficient search relative to other methods which repeatedly submit highly similar, incorrect solutions.
@@ 3 Motivation
Figure 3: Despite DeepSeek-Coder-V2-Lite-Base having significantly lower pass@1 than its instruct counterpart, we observe that this trend reverses as $k$ increases, suggesting that the instruct model has less diversity than its base model counterpart. We observe this trend for many, but not all, models and benchmarks, and provide the full data in Appendix H.
@@ 3.2 Backtranslation
To investigate the hypothesis whether the idea space, instantiated as solution sketches, is the right area of exploration, a natural question is whether LLMs can correctly implement a correct code solution given a correct sketch. Inspired by approaches to backtranslation in machine learning [35, 32, 16], we experiment with “backtranslating” passing code solutions back into idea space. First, we generate code solutions using GPT-4o to generate $1000$ attempts to solve the problem and filter out problems without any passing solutions. As we also do not have a dataset of correct solution sketches associated with each solution, we generate a candidate correct idea via backtranslation. We do this by feeding an LLM both the problem and code solution and asking the LLM to convert said solution into a natural language description of the solution. Additionally, we vary the detail of the backtranslated idea via instructions to the LLM in the prompt (e.g. ‘in $w$ words’). A full description of the prompts can be found in Appendix J.1, alongside several example backtranslated solutions of various lengths.
@@ 3.3 Conditioning on Idea Quality
In a follow-up experiment, we prompt an LLM to generate its own sketches to solve LiveCodeBench problems instead of providing it with golden ones via backtranslation. First, we generate $5$ ideas per problem using IdeaSearch, defined in Section 4.1.2. For each idea, we then sample $25$ candidate solutions and measure their pass rate. For this experiment, we filter out any problem that GPT-4o-mini solves with either a $100\%$ or a $0\%$ solve rate, since such problems are either too easy or too hard for the model and would not be informative for this experiment. We end with $75$ problems and $375$ sketches.
@@
Formally, for any problem $P_{i}$, we sample some sketch $I$ from some conditional distribution with probability mass $P(I|P_{i})$. The probability of solving $P_{i}$ is then $P(\mathrm{solve}|P_{i},I)$. We compare the solve-rate distribution, $P(\mathrm{solve}|P_{i},I)$ over all problems and all sketches versus the solve-rate distribution of $\sum_{I}P(\mathrm{solve}|P_{i},I)\cdot P(I|P_{i})=P(\mathrm{solve}|P_{i})$ over all problems.
@@
Our results confirm this to be the case. Figure 4(b) shows the distribution of solve rates across problems, both unconditionally (in red) and conditioned on each sketch (in blue). We notice that when grouping by sketches, the solve rates indeed become polarized towards $\{0,1\}$. This result has important implications for improving code generation, suggesting that a large portion of variance in performance may come from whether the model is able to generate a correct idea or not. Therefore, a natural path for improvement is to focus on the sketch generation step and search for correct sketches and observations in idea space before generating solution code.
@@ 4 Methods
We provide a description of the various methods of search we explore in our work. If additional background on competitive programming and related notation is desired, we provide more (optional) information in Appendix K.
4.1 Baselines
4.1.1 Repeated Sampling
We consider the basic prompting approach as a baseline, in which we use few-shot prompting by providing the LLM with a number of problem-solution pairs before asking it to solve the desired question [9]. A full example of the prompt is given in Appendix J.2. In code generation, the most common variant of search utilized is repeated sampling, where models are repeatedly sampled from until they generate an output that passes the test or the maximum number of samples is reached. Refer to the Related Work for more information (Section 2.2).
4.1.2 IdeaSearch
A natural extension of the Repeated Sampling approach discussed in Section 4.1.1 is to avoid prompting the LLM for the solution code immediately. This can be viewed as an application of the commonly used “chain-of-thought” prompting to programming problems [41], although we find that IdeaSearch shows non-negligible performance boosts over standard “chain-of-thought” prompting (see Appendix E).
In IdeaSearch, the LLM is given the problem $P$ and is asked to output a natural language solution $S$ of the problem. Then, a separate instance of the LLM is given $P$ and $S$, and tasked to follow the proposed solution $S$ to solve the problem $P$. The purpose of IdeaSearch is to isolate the effectiveness of having the correct “idea/sketch” for solving the problem. Empirically, we find that explicitly forcing the search algorithm to articulate an idea for solving the problem increases diversity. See Appendix J.3 for detailed prompts.
@@ 4.2.1 Prompting for Observations
We denote these observations as $O^{1}_{i}$, where, $i\in\{1,\dots,n_{1}\}$ due to the fact that they are first-order observations. Typically, $n_{1}$ is on the order of $3$ to $6$. The exact number depends on the LLM output. To use these observations to inspire future idea generation, we create all subsets with size at most $2$ of $S^{1}=\left\{O^{1}_{1},\dots,O^{1}_{n_{1}}\right\}$. Each of these subsets is a combination of observations, and for clarity we denote each subset as $C^{1}_{i},i\in\{1,\dots,l_{1}\}$, where $l_{1}=1+n_{1}+\binom{n_{1}}{2}$.
@@ 4.2.2 Deriving New Observations
The same procedure as Section 4.2.1 is used to create all subsets $C^{2}_{i,j}$, for all $i\in\{1,\dots,l_{1}\}$. This process may be arbitrarily repeated, but we truncate the tree at depth $2$ for computational constraints.
@@ 4.2.3 Observations to Code
These natural language solutions are then translated into pseudocode, which are subsequently translated into actual Python code. We take a more granular approach to reduce the translation error (which may cause the model to revert to its original mode, disregarding the reasoned-through observations). We provide all prompts for all sections in Appendix J.4.
@@ 5 Experimental Results
Model Benchmark Pass@1 Pass@200 IdeaSearch@200 PlanSearch@200
GPT-4o-mini LiveCodeBench 39.0 53.3 59.4 64.9
GPT-4o LiveCodeBench 41.3 60.6 70.4 73.0
DeepSeek-Coder-V2 LiveCodeBench 41.4 53.2 65.9 70.3
Claude-Sonnet-3.5 LiveCodeBench 40.3 55.6 70.2 77.0
GPT-4o-mini HumanEval+ 83.7 95.0 97.5 98.2
GPT-4o HumanEval+ 86.4 98.2 97.6 99.5
DeepSeek-Coder-V2 HumanEval+ 82.8 91.4 97.2 99.3
Claude-Sonnet-3.5 HumanEval+ 81.6 88.9 95.6 98.5
GPT-4o-mini MBPP+ 73.5 83.8 87.3 91.0
GPT-4o MBPP+ 77.2 87.4 89.3 92.2
DeepSeek-Coder-V2 MBPP+ 76.3 81.9 89.1 92.6
Claude-Sonnet-3.5 MBPP+ 77.1 83.0 87.8 93.7
Table 1: We find that PlanSearch and IdeaSearch improve upon search baselines across all models, with PlanSearch achieving the best results across all models and benchmarks considered. Notably, using PlanSearch on top of Claude 3.5 Sonnet [2] has a pass@200 of $77.0$ on LiveCodeBench, which is nearly double the performance of the top model without using search ($41.4$). We highly encourage readers to check Appendix A for complete results and pass@k curves.
Figure 5: Performance of all models and methods on LiveCodeBench with public test filtering. The idea of public test filtering is to shift pass@k curves leftward (i.e., bringing high $k$ to low $k$), so we plot curves in detail over $k\in\{1,\dots,20\}$. Dotted lines are provided for reference of the base method pass@k before filtering. Even at 10 completions, PlanSearch outperforms filtered Repeated Sampling by a flat 30 to 40%. Again, full pass@k plots are included in their entirety in Appendix A.

### v2

@@ Abstract
While scaling training compute has led to remarkable improvements in large language models (LLMs), scaling inference compute has not yet yielded analogous gains. We hypothesize that a core missing component is a lack of diverse LLM outputs, leading to inefficient search due to models repeatedly sampling highly similar, yet incorrect generations. We empirically demonstrate that this lack of diversity can be mitigated by searching over candidate plans for solving a problem in natural language. Based on this insight, we propose PlanSearch, a novel search algorithm which shows strong results across HumanEval+, MBPP+, and LiveCodeBench (a contamination-free benchmark for competitive coding). PlanSearch generates a diverse set of observations about the problem and uses these observations to construct plans for solving the problem. By searching over plans in natural language rather than directly over code solutions, PlanSearch explores a significantly more diverse range of potential solutions compared to baseline search methods. Using PlanSearch on top of Claude 3.5 Sonnet achieves a pass@200 of 77.0% on LiveCodeBench, outperforming both the best pass-rate achieved without any search (pass@1 = 41.4%) and using standard repeated sampling on top of existing non-search models (pass@200 = 60.6%). Finally, we show that, across all models, search algorithms, and benchmarks analyzed, we can accurately predict performance gains from search as a function of the diversity over generated ideas. Code can be found at https://github.com/scaleapi/plansearch.
@@ 1 Introduction
The bitter lesson [45] famously posits that two forms of scaling trump everything else: learning and search. While recent advances in large language models (LLMs) have shown that learning is extremely effective, search has not yet proven its value for LLMs, despite its success with classical machine learning techniques [10, 42, 43, 7, 8, 4, 18].
@@
Furthermore, the lack of diversity is particularly harmful for search algorithms. In the most egregious of cases with little to no diversity, such as greedy decoding, repeated sampling from the model returns highly similar programs, resulting in minimal gain from additional inference-time compute. This diversity problem is also not reflected in many public leaderboards (e.g. LMSYS Chatbot Arena [15], LiveCodeBench [24], OpenLLMLeaderboard [1]), which often report only the pass rate from a single sample of the model, ignoring an entire dimension along which to compare models. While the performance of one sample is the primary metric of relevance for applications such as chatbots, as users typically are sensitive to latency, this single scalar is insufficient to fully capture the quality of a model when it is allowed to use more inference-time compute.
@@
In this paper, we explore several directions for improving the diversity of LLMs at inference time. We hypothesize that the right axis of diversity to search over is the natural language conceptual/idea space, and we validate our hypothesis across several experiments. First, we show that models can produce the correct final program when fed correct solution sketches, where these sketches have been “backtranslated” from passing solution code into sketches in idea space (Section 3.2). Second, we show that when models are asked to generate their own ideas before implementing them on LiveCodeBench (IdeaSearch), their accuracy conditioned on a particular sketch trends towards either 0% or 100%, suggesting that most of the variance in passing a particular problem is captured by whether the sketch is correct rather than any other factor. These two experiments suggest a natural method to improving LLM search for code generation: by searching for the correct idea to implement.
Guided by this principle of maximizing exploration of ideas, we propose PlanSearch. In contrast to many existing search methods that search over individual tokens [53, 56], lines of code [26], or even entire programs [27], PlanSearch searches over possible plans for solving the problem at hand, where a plan is defined as a collection of high level observations and sketches helpful to solve a particular problem (Figure 2). To generate novel plans, PlanSearch generates a number of observations about the problem, before combining these observations into a candidate plan for solving the problem. This is done for every possible subset of the generated observations to maximally encourage exploration in idea space, before the codes are eventually all translated into a final code solution (Section 4.3). We find that searching over plans outperforms both standard repeated sampling and directly searching over ideas (IdeaSearch, introduced in Section 4.2) in terms of effectively using compute at inference time.
Applying PlanSearch on top of Claude 3.5 Sonnet achieves a pass@200 of 77.0% on LiveCodeBench, outperforming both the best score achieved without search (pass@1 = 41.4%) and the standard best-of-n sampling score on non-search models (pass@200 = 60.6%). Furthermore, consistent with recent findings on the effectiveness of search on top of small models [13, 6, 5, 47], running PlanSearch on top of a small model (GPT-4o-mini) outperforms larger models not augmented with search after merely 4 attempts. Evaluations of PlanSearch across two other coding benchmarks, HumanEval+ and MBPP+ [28], suggest similar improvements.
@@ 2.1 Search in Classical AI
Classical search algorithms like breadth-first search, depth-first search, and A* search have been widely used for pathfinding, planning, and optimization [37]. More advanced search techniques like Monte Carlo Tree Search (MCTS) have achieved remarkable success in domains like game playing, enabling superhuman performance in Go [42, 43], Poker [7, 8] and Diplomacy [18]. More recently, scaling laws have been found for the performance of AI systems in board games, where ELO improves logarithmically with the amount of compute spent at inference [25].
@@ 2.2 Search with Language Models
Regarding searching over plans in natural language, several approaches have also proposed generalizing chain-of-thought [46] reasoning into a search-like process, such as Tree of Thoughts [48] and Reasoning via Planning [22]. However, prior methods have largely demonstrated effectiveness on somewhat contrived problems designed to highlight the power of search, such as the game of 24, or classic planning benchmarks such as Blocksworld [29], where both benchmarks are easier to solve by explicitly considering many options, and where the ‘steps’ over which to search over are fairly obvious. By contrast, most real-world planning is used to assist in domains that are complex enough to benefit from, but not require, the additional exploration of plans. We demonstrate that PlanSearch, which plans in natural language, outperforms baseline search methods in one such domain: code generation. Moreover, our analysis reveals the underlying reason that such search is effective: it increases the diversity of the generated ideas, allowing more efficient search relative to other methods which repeatedly submit highly similar, incorrect solutions. This is consistent with prior work suggesting the importance of diversity in natural language generation [23, 55]. Other directions for search include decomposing programs down into smaller parts before solving each one individually [51, 57, 20].
PlanSearch is also distinct from other methods which explicitly train a model to search or on reasoning traces sampled from the model [50, 54, 52, 32] in that PlanSearch induces diversity at inference-time and converts an LLM API not designed for search into one that is capable of showing strong gains from search. Separately, there is a large family of work in the agent space, in which outputs from terminal commands or other tools are fed back into the model before the agent is queried for the next step [41, 49, 38, 12].
@@ 3 Motivation
Figure 3: Despite DeepSeek-Coder-V2-Lite-Base having significantly lower pass@1 than its instruct counterpart, we observe that this trend reverses as $k$ increases, suggesting that the instruct model has less diversity than its base model counterpart. We observe this trend for many, but not all, models and benchmarks, and provide the full data in Appendix I.
@@ 3.2 Backtranslation
To investigate the hypothesis whether the idea space, instantiated as solution sketches, is the right area of exploration, a natural question is whether LLMs can correctly implement a correct code solution given a correct sketch. Inspired by approaches to backtranslation in machine learning [39, 35, 17], we experiment with “backtranslating” passing code solutions back into idea space. First, we generate code solutions using GPT-4o to generate $1000$ attempts to solve the problem and filter out problems without any passing solutions. As we also do not have a dataset of correct solution sketches associated with each solution, we generate a candidate correct idea via backtranslation. We do this by feeding an LLM both the problem and code solution and asking the LLM to convert said solution into a natural language description of the solution. Additionally, we vary the detail of the backtranslated idea via instructions to the LLM in the prompt (e.g. ‘in $w$ words’). A full description of the prompts can be found in Appendix M.1, alongside several example backtranslated solutions of various lengths.
@@ 3.3 Conditioning on Idea Quality
In a follow-up experiment, we prompt an LLM to generate its own sketches to solve LiveCodeBench problems instead of providing it with golden ones via backtranslation. First, we generate $5$ ideas per problem using IdeaSearch, defined in Section 4.2. For each idea, we then sample $25$ candidate solutions and measure their pass rate. For this experiment, we filter out any problem that GPT-4o-mini solves with either a $100\%$ or a $0\%$ solve rate, since such problems are either too easy or too hard for the model and would not be informative for this experiment. We end with $75$ problems and $375$ sketches.
@@
Formally, for a problem $P_{i}$, IdeaSearch samples a sketch $I$ from some conditional distribution with probability mass $P(I|P_{i})$. The probability of solving $P_{i}$ is then $P(\mathrm{solve}|P_{i},I)$. In this experiment, we compare the solve-rate distribution $P(\mathrm{solve}|P_{i},I)$ over all problems and all sketches versus the solve-rate distribution of $\sum_{I}P(\mathrm{solve}|P_{i},I)\cdot P(I|P_{i})=P(\mathrm{solve}|P_{i})$ over all problems.
@@
Our results confirm this to be the case. Figure 4(b) shows the distribution of solve rates across problems, both unconditionally (in red) and conditioned on each sketch (in blue). We notice that when grouping by sketches, the solve rates indeed become polarized towards $\{0,1\}$. This result has important implications for improving code generation, suggesting that a large portion of variance in performance can be explained by whether the model is able to generate a correct idea or not. Therefore, a natural path for improvement is to focus on the sketch generation step and search for correct sketches and observations in idea space before generating solution code.
@@ 4 Methods
We provide a description of the various methods of search we explore in our work. If additional background on competitive programming and related notation is desired, we provide more (optional) information in Appendix P.
4.1 Repeated Sampling
We consider the basic prompting approach as a baseline, in which we use few-shot prompting by providing the LLM with a number of problem-solution pairs before asking it to solve the desired question [9]. A full example of the prompt is given in Appendix M.2. In code generation, the most common variant of search utilized is repeated sampling, where models are repeatedly sampled from until they generate an output that passes the test or the maximum number of samples is reached. Refer to the Related Work for more information (Section 2.2).
4.2 IdeaSearch
A natural extension of the Repeated Sampling approach discussed in Section 4.1 is to avoid prompting the LLM for the solution code immediately. This can be viewed as an application of the commonly used “chain-of-thought” prompting to programming problems [46], although we find that IdeaSearch shows non-negligible performance boosts over standard “chain-of-thought” prompting (see Appendix E).
In IdeaSearch, the LLM is given the problem $P$ and is asked to output a natural language solution $S$ of the problem. Then, a separate instance of the LLM is given $P$ and $S$, and tasked to follow the proposed solution $S$ to solve the problem $P$. The purpose of IdeaSearch is to isolate the effectiveness of having the correct “idea/sketch” for solving the problem. Empirically, we find that explicitly forcing the search algorithm to articulate an idea for solving the problem increases diversity. See Appendix M.3 for detailed prompts.
@@ 4.3 PlanSearch
Ablations for many of the choices in the subsequent description of PlanSearch can be found in Appendix H.
@@ 4.3.1 Prompting for Observations
We denote these observations as $O^{1}_{i}$, where, $i\in\{1,\dots,n_{1}\}$ due to the fact that they are first-order observations. Typically, $n_{1}$ is on the order of $3$ to $6$. The exact number depends on the LLM output. To use these observations to inspire future idea generation, we create all subsets with size at most $S=2$ of $s^{1}=\left\{O^{1}_{1},\dots,O^{1}_{n_{1}}\right\}$. Each of these subsets is a combination of observations, and for clarity we denote each subset as $C^{1}_{i},i\in\{1,\dots,l_{1}\}$, where $l_{1}=1+n_{1}+\binom{n_{1}}{2}$.
@@ 4.3.2 Deriving New Observations
The same procedure as Section 4.3.1 is used to create all subsets $C^{2}_{i,j}$, for all $i\in\{1,\dots,l_{1}\}$. This process may be arbitrarily repeated, but we truncate the tree at depth $L=2$ for computational constraints.
@@ 4.3.3 Observations to Code
These natural language solutions are then translated into pseudocode, which are subsequently translated into actual Python code. We take a more granular approach to reduce the translation error (which may cause the model to revert to its original mode, disregarding the reasoned-through observations). We provide all prompts for all sections in Appendix M.4.
@@ 5 Experimental Results
Model Eval Pass@1 Pass@200 IS@200 (ours) PS@200 (ours)
GPT-4o-mini LCB 39.0 53.3 59.4 64.9
GPT-4o LCB 41.3 60.6 70.4 73.0
DeepSeek-Coder-V2 LCB 41.4 53.2 65.9 70.3
Claude-Sonnet-3.5 LCB 40.3 55.6 70.2 77.0
GPT-4o-mini HE+ 83.7 95.0 97.5 98.2
GPT-4o HE+ 86.4 98.2 97.6 99.5
DeepSeek-Coder-V2 HE+ 82.8 91.4 97.2 99.3
Claude-Sonnet-3.5 HE+ 81.6 88.9 95.6 98.5
GPT-4o-mini M+ 73.5 83.8 87.3 91.0
GPT-4o M+ 77.2 87.4 89.3 92.2
DeepSeek-Coder-V2 M+ 76.3 81.9 89.1 92.6
Claude-Sonnet-3.5 M+ 77.1 83.0 87.8 93.7
o1-mini (search model) LCB 69.5 90.8 91.2 91.3
Table 1: LCB, HE+, M+ short for LiveCodeBench, HumanEval+, and MBPP+, respectively. IS short for IdeaSearch and PS short for PlanSearch. We find that PlanSearch and IdeaSearch improve upon search baselines across all models, with PlanSearch achieving the best results across all models and benchmarks considered. Notably, using PlanSearch on top of Claude 3.5 Sonnet [2] has a pass@200 of $77.0$ on LiveCodeBench, which is nearly double the performance of the top model without using search ($41.4$). PlanSearch also outperforms basic pass@200 on o1-mini for LiveCodeBench, though since o1-mini already uses inference-time compute, the gap is much smaller than compared to non-search models. The full pass@k curves are included in Appendix A.
Figure 5: Performance of all models and methods on LiveCodeBench with public test filtering. The purpose of filtering is to shift pass@k curves leftward (i.e., bringing performance high $k$ to low $k$), so we plot curves in detail over $k\in\{1,\dots,20\}$. Dotted lines are provided for reference of the base method pass@k before filtering. Even at 10 completions, PlanSearch outperforms filtered Repeated Sampling by a flat 30 to 40%. Again, full pass@k plots are included in their entirety in Appendix A.